THE WOODBURY VEGETABLE GARDEN IS IN THE SHAPE OF A RECTANGLE
SURROUNDED BY 330 FEET OF FENCE THE LENGTH OF THE GARDEN IS 75 FEET MORE
THAN TWICE THE WIDTH OF THE GARDEN
FIND THE DIMENSIONS OF THIS GARDEN

Answer by waynest(281)

length (L) = 2w + 75
width = w
perimeter = 330
p = 2w + 2L
2w + 2(2w + 75) = 330
2w + 4w + 150 = 330
6w + 150 - 150 = 330 - 150
6w = 180
6w/6 = 180/6
w = 30
check:
2(30) + 2(2(30) + 75) = 330
60 + 2(60 + 75) = 330
60 + 2(135) = 330
60 + 270 = 330
330 = 330